OpenStudy (anonymous):

The table below shows the fees to check bags on a certain airline. If x represents the total number of bags a person checks and each person checks more than two bags, the equation below can be used to determine C, the total fee for all the bags. 20 + 25 + 50(x - 2) = C Arianna paid a total of $145 to check all her bags on a flight. How many bags did she check in all?

OpenStudy (ivettef365):

substitute the C for 145, and solve for x, so you have 20 + 25 + 50(x-2) = 145

OpenStudy (anonymous):

20+25+50x-100=145

OpenStudy (ivettef365):

good, now combine like terms

OpenStudy (anonymous):

45-55=145

OpenStudy (ivettef365):

no, what you do is add the numbers so you have 20 + 25 = 45 -100 = -55 so you have 50x -55 = 145

OpenStudy (ivettef365):

now add 55 on both sides ?

OpenStudy (anonymous):

4?

OpenStudy (ivettef365):

good job

OpenStudy (anonymous):

Thank you. & could you go look at the other problem I just posted, I can not figure that one for nothing.

OpenStudy (ivettef365):

sure